# CryoFastAR: Fast Cryo-EM Ab initio Reconstruction Made Easy

### v1

@@ Abstract
Pose estimation from unordered images is fundamental for 3D reconstruction, robotics, and scientific imaging. Recent geometric foundation models, such as DUSt3R, enable end-to-end dense 3D reconstruction but remain underexplored in scientific imaging fields like cryo-electron microscopy (cryo-EM) for near-atomic protein reconstruction. In cryo-EM, pose estimation and 3D reconstruction from unordered particle images still depend on time-consuming iterative optimization, primarily due to challenges such as low signal-to-noise ratios (SNR) and distortions from the contrast transfer function (CTF). We introduce CryoFastAR, the first geometric foundation model that can directly predict poses from Cryo-EM noisy images for Fast ab initio Reconstruction. By integrating multi-view features and training on large-scale simulated cryo-EM data with realistic noise and CTF modulations, CryoFastAR enhances pose estimation accuracy and generalization. To enhance training stability, we propose a progressive training strategy that first allows the model to extract essential features under simpler conditions before gradually increasing difficulty to improve robustness. Experiments show that CryoFastAR achieves comparable quality while significantly accelerating inference over traditional iterative approaches on both synthetic and real datasets. We will release our code, models, and datasets to stimulate further research.
@@ 1 Introduction
Recent geometric foundation models, such as DUSt3R [47] are transforming this paradigm toward a generalized per-image inference strategy, enabling end-to-end reconstruction from unposed images, significantly improving scalability for large-scale 3D reconstruction tasks. Nevertheless, these advancements remain largely unexplored in scientific imaging, particularly in cryo-electron microscopy (cryo-EM), where accurate pose estimation is fundamental to high-resolution protein reconstruction. In cryo-EM, jointly estimating poses and reconstructing 3D structures from hundreds of thousands of unordered particle images, a process known as ab initio reconstruction, still relies on per-scene iterative optimization. This is primarily due to several unique challenges, including extremely low signal-to-noise ratios (SNR)[42] and contrast transfer function (CTF) distortions[31].
@@
To address this, we present CryoFastAR, the first geometric foundation model that can directly predict poses from Cryo-EM unordered noisy images for Fast Ab initio Reconstruction. Our method employs a vision transformer-based architecture to directly estimate relative poses in a global coordinate system by efficiently extracting and integrating multi-view image features. Instead of explicitly regressing conventional 5D pose parameters, CryoFastAR predicts a dense Fourier planar map encoding each image’s pose relative to a reference view, enabling more flexible optimization. This representation can be readily converted into conventional 5D pose parameters for seamless integration with existing reconstruction algorithms. Additionally, by training on large-scale, realistically simulated cryo-EM data, CryoFastAR achieves robust pose estimation without requiring precomputed contrast transfer function (CTF) parameters, simplifying the reconstruction pipeline and enhancing generalization capability.
Our model is trained in a fully supervised manner to predict relative poses from multi-view cryo-EM images. We curate a large-scale cryo-EM simulation dataset consisting of over 110,000 protein structures derived from publicly available PDB models [3], with each structure associated with 100 simulated images and corresponding ground-truth poses. The diversity of these structures provides extensive geometric variation essential for robust training. To enhance generalization to real data, we apply realistic data augmentations, including random noise addition, CTF simulation, and other perturbations representative of experimental conditions. Moreover, we adopt a progressive training strategy, initially training on clean images with fewer views, then gradually increasing noise levels, CTF distortions, and the number of input views. This staged approach enables the model to first learn essential structural features and subsequently adapt to more challenging data distributions. Finally, we fine-tune the model using a small set of real cryo-EM images of complex proteins, incorporating the maximum number of views to further improve robustness.
We extensively evaluate CryoFastAR on both unseen synthetic and experimental cryo-EM datasets. On synthetic datasets, CryoFastAR achieves state-of-the-art results in terms of 2D in-plane shift estimation and overall reconstruction quality. With a rapid pose refinement step in CryoSPARC [30], our method achieves superior performance at substantially lower computational cost, accelerating reconstruction by over an order of magnitude (10×) compared to existing methods. On real datasets, CryoFastAR demonstrates reconstruction results comparable to current state-of-the-art methods, yet significantly reduces computational time. Our code, models, and datasets will be publicly available to facilitate future research.
@@ 2 Related Work
Neural Representations for Cryo-EM Reconstruction. Neural representations have recently emerged as promising tools for addressing continuous structural heterogeneity in cryo-EM reconstruction [53, 19, 20, 37, 22, 25, 14, 23, 27]. CryoDRGN [53] introduced a VAE architecture that encodes conformational states from images and decodes them through a coordinate-based MLP representing the 3D Fourier volume. However, this approach still relies on traditional pipelines for pose estimation and assumes a static structure. More recent neural approaches [54, 19, 38] attempt to jointly optimize pose estimation and neural volume reconstruction. CryoDRGN2 [54] and DRGN-AI [21] integrates iterative pose optimization with neural reconstruction, while CryoAI [19] employs amortized inference for direct pose estimation, and CryoSPIN [38] further improves this with a semi-amortized strategy. Nevertheless, all these methods remain limited by computationally intensive per-scene optimization, often resulting in slow convergence and susceptibility to local minima. In contrast, our approach directly predicts Fourier planar maps in a fully feed-forward manner from multi-view particle images, enabling more efficient and accurate reconstructions.
@@ 3.1 Image Formation Model
Cryo-EM image formation can be modeled by assuming each observed particle image is generated from an unknown underlying 3D electron density map $V:\mathbb{R}^{3}\mapsto\mathbb{R}$. Each particle undergoes an unknown rotation $R\in\text{SO}(3)$ and an in-plane translation $\mathbf{t}=(t_{x},t_{y})^{\top}\in\mathbb{R}^{2}$, shifting the projected image by $t_{x}$ pixels along the $x$ axis and $t_{y}$ pixels along the $y$ axis, respectively. The projection operator $\mathcal{P}_{R,\mathbf{t}}$ integrates the density along the optical axis (the $z$-axis), formally defined as:
@@ 3.2 Fourier Slicing Theorem
where $\text{CTF}_{i}$ is the microscope’s contrast transfer function, which is the 2D transform of its corresponding $\text{PSF}_{i}$, and $\hat{\epsilon}_{i}$ denotes additive Gaussian noise in the frequency domain. The operator $\odot$ represents element-wise multiplication.
@@ 3.3 Homogeneous Reconstruction
where the set $\mathcal{K}(\boldsymbol{\omega})=\left\{\mathbf{k}_{i}=h^{-1}\left(R_{i}^{\top}\boldsymbol{\omega}\right)\ \big{|}\ \mathbf{r}_{i}^{\top}\boldsymbol{\omega}=0\right\}$ collects the relevant Fourier coordinates. Here, $\mathbf{r}_{i}$ denotes the third column vector of $R_{i}$, and the inverse homogeneous operator $h^{-1}:\mathbb{R}^{3}\mapsto\mathbb{R}^{2}$ is defined as $h^{-1}(\boldsymbol{\omega})=\left(\omega_{x},\omega_{y}\right)^{\top}.$ Finally, the inverse 3D Fourier transform is applied to obtain the real-space volume: $V=\mathcal{F}^{-1}_{\text{3D}}(\hat{V}).$
@@ 4 Methods
As shown in Figure 2, CryoFastAR employs a vision transformer-based architecture that efficiently encodes particle image features and decodes them into multi-view Fourier planar maps. Specifically, a shared ViT encoder extracts patch-level features, which are then refined through stacked view integration blocks to produce informative representations. These representations are subsequently decoded to Fourier planar maps encoding relative poses (Section 4.1). To effectively train our model on large-scale datasets with high noise levels, we adopt a progressive training strategy, gradually increasing data complexity and noise levels (Section 4.2). During inference, the predicted Fourier planar maps are regressed to explicit 5D poses, facilitating fast ab initio protein reconstruction (Section 4.3).
@@
A fundamental challenge in protein structure recovery is accurately predicting the orientation of a slicing plane in the canonical Fourier space. To address this, we introduce a novel representation termed the Fourier planar map, which characterizes a shifted slice in the Fourier space as a dense map corresponding to a 5D pose $(R,\mathbf{t})$:
@@
where $X^{0}$ represents a uniformly sampled 2D grid on the plane $z=0$, spanning the range $[-1,1]^{2}$. Our goal is to develop a neural network that directly takes a set of CryoEM images as input and predicts the corresponding Fourier planar map.
@@ 4.1 Model Architecture
Directly applying self-attention across all views introduces quadratic memory and computational complexity, becoming impractical for scenarios involving dozens or hundreds of particle images. To address this, we propose an efficient cross-attention-based approach whose complexity scales linearly with the number of views. Specifically, each decoder layer consists of two key modules: 1) a view integration block, which aggregates target views’ features into a primary reference view (we only choose the first view) via cross-attention; and 2) a view updating block, which further refines auxiliary views conditioned on the updated primary features.
@@
Finally, the updated feature representation of each view is decoded into a pixel-wise Fourier planar map accompanied by an auxiliary confidence map, expressed within the reference (first) view’s 3D Fourier coordinate frame. Specifically, two separate downstream heads perform this decoding: one head predicts the primary (reference) view’s planar map and the other predicts planar maps for all target views. Formally expressed as:
@@ 4.2 Training
Directly training the model end-to-end on high-noise cryo-EM images poses significant convergence challenges due to the significant differences between cryo-EM data and typical computer vision datasets. To overcome this, we adopt a progressive training approach consisting of three stages. First, we pre-train our model on clean projection images using only two views per particle, facilitating rapid convergence on simplified conditions. Next, we progressively expand training to our full-scale dataset, gradually increasing input views from 2 to 16 and simultaneously reducing the SNR from 10.0 to 0.1, with added CTF distortions. Finally, to mitigate underfitting on experimental data, we fine-tune our model on a small set of real cryo-EM images. This progressive training strategy ensures stable convergence, enabling the model to robustly extract and integrate multi-view features for accurate ab initio reconstruction.
@@ 4.3 Inference
After training, our model generalizes to an arbitrary number of input views. Although we initially trained it with 32 views, we found that increasing the number of input views improves performance as our model successfully learned to integrate multi-view information. In our experiments, we utilize 128 views during inference, setting the first view as the reference to predict the Fourier planar maps of the remaining views in many batches.
@@ 5 Experiments
To train our model, we construct the first large-scale simulated cryo-EM dataset containing 113,600 atomic structures of protein complexes curated from the Protein Data Bank (PDB). For each atomic structure, we first generate a density volume at a resolution of $128^{3}$ using EMAN2 [41], with voxel sizes automatically determined (ranging from 1Å to 10Å) to fully enclose each structure. From each density volume, we uniformly sample orientations to produce 100 clean 2D projection images at $128\times 128$ resolution. To simulate realistic imaging conditions, we further apply data augmentations, including additive random noise, random in-plane 2D shifts, and CTF modulation. The CTF parameters are sampled from 152,385 CTF parameters from a large-scale real micrograph dataset [39]. For real-world training, we utilize four experimental datasets from CryoDRGN-EMPIAR [53], each containing accurate pre-computed particle poses. For each experimental dataset, we randomly select 30,000 images for training and exclude them from evaluation.
@@ 5.1 Comparison
We compare our approach against three ab initio reconstruction baselines: (1) CryoSPARC, a widely-used software utilizing iterative expectation-maximization for pose estimation; (2) CryoSPIN, a neural method directly predicting image poses from single-view inputs; and (3) CryoDRGN2, a hybrid method alternating neural reconstruction and iterative pose search. Notably, unlike these methods, our approach does not require pre-computed CTF parameters for pose estimation. All baseline methods rely on provided CTF parameters during iterative pose refinement or per-scene optimization. Since the official implementation of CryoSPIN does not support shift estimation, we disable random shifts when training it on simulated data and exclude it from comparisons on experimental data. After initial ab initio reconstruction, we further refine our model’s predicted poses using CryoSPARC’s homogeneous refinement to evaluate the quality of our initialization compared to CryoSPARC’s own initialization. Specifically, given our estimated poses, we run CryoSPARC’s reconstruction only followed by the local refinement. For CryoSPARC, we continue to run homogeneous refinement followed by a local refinement.
@@
We compare our method with baseline approaches both qualitatively and quantitatively. As shown in Figure 3, our approach produces reconstructions with superior structural completeness and finer detail preservation. Quantitative results reported in Table 1 further confirm that our method outperforms baselines in pose estimation accuracy and reconstruction quality, achieving significantly lower rotation and translation errors. Remarkably, our method accelerates the reconstruction process by over an order of magnitude (10×), while maintaining or surpassing the quality of all baselines. After refinement, our approach consistently achieves the best overall performance, demonstrating CryoFastAR’s strong initialization quality, computational efficiency, and robustness in challenging reconstruction scenarios.
Experimental results. Table 2 summarizes the quantitative results on three experimental cryo-EM datasets including RAG, 50S ribosome, and pre-catalytic Spliceosome [26]. Overall, our approach achieves competitive reconstruction accuracy compared to baseline methods while significantly reducing the computational cost. Specifically, our method achieves approximately a $3\times3$ speed-up compared to CryoSPARC and over a 33.21×33.21$\times$33.21 × speed-up relative to CryoDRGN2. Notably, our method performs particularly well on the RAG and spliceosome datasets, attaining comparable or better pose estimation accuracy relative to the baselines. However, on the 50S ribosome dataset, our method, although predicting poses in generally correct orientations, yields substantially lower accuracy than the baselines. We hypothesize that this discrepancy arises from the intrinsic structural flexibility and complexity associated with membrane proteins like 50S, posing greater challenges for our feed-forward approach trained primarily on simulated globular structures. As demonstrated in Figure 4, we provide visual comparisons of reconstructions on the experimental spliceosome dataset. These visualizations confirm our quantitative comparisons, demonstrating that our refined reconstruction exhibits improved structural fidelity and better captures the detailed molecular features compared to baseline methods. Other baselines may fail due to lacking the robustness of their heterogeneity without direct supervision. The FSC curves further support this observation, showing improved reconstruction resolution after refinement.

### v2

@@ Abstract
Pose estimation from unordered images is fundamental for 3D reconstruction, robotics, and scientific imaging. Recent geometric foundation models, such as DUSt3R, enable end-to-end dense 3D reconstruction but remain underexplored in scientific imaging fields like cryo-electron microscopy (cryo-EM) for near-atomic protein reconstruction. In cryo-EM, pose estimation and 3D reconstruction from unordered particle images still depend on time-consuming iterative optimization, primarily due to challenges such as low signal-to-noise ratios (SNR) and distortions from the contrast transfer function (CTF). We introduce CryoFastAR, the first geometric foundation model that can directly predict poses from Cryo-EM noisy images for fast, feed-forward ab initio reconstruction. By integrating multi-view features and training on large-scale simulated cryo-EM data with realistic noise and CTF modulations, CryoFastAR enhances pose estimation accuracy and generalization. To enhance training stability, we propose a progressive training strategy that first allows the model to extract essential features under simpler conditions before gradually increasing difficulty to improve robustness. Experiments show that CryoFastAR achieves comparable quality while significantly accelerating inference over traditional iterative approaches on both synthetic and real datasets.
@@ 1 Introduction
Recent geometric foundation models, such as DUSt3R [47] are transforming this paradigm toward a generalized per-image inference strategy, enabling end-to-end reconstruction from unposed images, significantly improving performance for large-scale 3D reconstruction tasks. Nevertheless, these advancements remain largely unexplored in scientific imaging, particularly in cryo-electron microscopy (cryo-EM), where accurate pose estimation is fundamental to high-resolution protein reconstruction. In cryo-EM, jointly estimating poses and reconstructing 3D structures from hundreds of thousands of unordered particle images, a process known as ab initio reconstruction, still relies on per-scene iterative optimization. This is primarily due to several unique challenges in cryo-EM, including extremely low signal-to-noise ratios (SNR)[43] and contrast transfer function (CTF) distortions[32].
@@
To address this, we present CryoFastAR, the first geometric foundation model that can directly predict poses from Cryo-EM unordered noisy images for Fast Ab initio Reconstruction. Our method employs a vision transformer-based architecture to directly estimate relative poses in a global coordinate system by efficiently extracting and integrating multi-view image features. Instead of explicitly regressing conventional 5D pose parameters, CryoFastAR predicts a dense Fourier planar map to encode each image’s pose relative to a reference view, enabling more flexible optimization. This representation can be readily converted into conventional 5D pose parameters for seamless integration with existing reconstruction algorithms. Additionally, by training on large-scale, realistically simulated cryo-EM data, CryoFastAR achieves robust pose estimation without requiring precomputed contrast transfer function (CTF) parameters, simplifying the reconstruction pipeline and enhancing generalization capability.
Our model is trained in a fully supervised manner to predict relative poses from multi-view cryo-EM images. We curate a large-scale cryo-EM simulation dataset consisting of over 110,000 protein structures derived from publicly available PDB models [3], with each structure associated with 100 simulated images and corresponding ground-truth poses. The diversity of these structures provides extensive geometric variation essential for robust training. To enhance generalization to real data, we apply realistic data augmentations, including random noise addition, CTF simulation, and other perturbations representative of experimental conditions. Moreover, we adopt a progressive training strategy, initially training on clean images with fewer views, then gradually increasing noise levels, CTF distortions, and the number of input views. This staged approach enables the model to first learn essential structural features and subsequently adapt to more challenging data distributions. Finally, we fine-tune the model using a small set of real cryo-EM images of complex proteins, incorporating the maximum number of views to further improve performance on real datasets.
We extensively evaluate CryoFastAR on both unseen synthetic and experimental cryo-EM datasets. On synthetic datasets, CryoFastAR achieves state-of-the-art results in terms of 2D in-plane shift estimation and overall reconstruction quality. With a rapid pose refinement step in CryoSPARC [31], our method achieves superior performance at substantially lower computational cost, accelerating reconstruction by over an order of magnitude (10×) compared to existing methods. On real datasets, CryoFastAR demonstrates reconstruction results comparable to current state-of-the-art methods, yet significantly reduces computational time.
@@ 2 Related Work
Neural Representations for Cryo-EM Reconstruction. Neural representations have recently emerged as promising tools for addressing continuous structural heterogeneity in cryo-EM reconstruction [53, 20, 21, 38, 23, 26, 15, 24, 28]. CryoDRGN [53] introduced a variational autoencoder (VAE) architecture that encodes conformational states from images and decodes them through a coordinate-based multi-layer perceptrons (MLP) representing the 3D Fourier volume. However, this approach still relies on traditional pipelines for pose estimation and assumes a static structure. More recent neural approaches [54, 20, 39] attempt to jointly optimize pose estimation and neural volume reconstruction. CryoDRGN2 [54] and DRGN-AI [22] integrates iterative pose optimization with neural reconstruction, while CryoAI [20] employs amortized inference for direct pose estimation, and CryoSPIN [39] further improves this with a semi-amortized strategy. Nevertheless, all these methods remain limited by computationally intensive per-scene optimization, often resulting in slow convergence and susceptibility to local minima. In contrast, our approach directly predicts Fourier planar maps in a fully feed-forward manner from multi-view particle images, enabling more efficient and accurate reconstructions.
@@ 3.1 Image Formation Model
Cryo-EM image formation can be modeled by assuming each observed particle image is generated from an unknown underlying 3D electron density map $V:\mathbb{R}^{3}\mapsto\mathbb{R}$. Each particle undergoes an unknown rotation $R\in\text{SO}(3)$ and an in-plane translation $\mathbf{t}=(t_{x},t_{y})^{\top}\in\mathbb{R}^{2}$ in the camera coordinates, shifting the projected image by $t_{x}$ pixels along the $x$ axis and $t_{y}$ pixels along the $y$ axis, respectively. The projection operator $\mathcal{P}_{R,\mathbf{t}}$ integrates the density along the optical axis (the $z$-axis), formally defined as:
@@ 3.2 Fourier Slicing Theorem
where $\text{CTF}_{i}$ is the microscope’s contrast transfer function, which is the 2D Fourier transform of its corresponding $\text{PSF}_{i}$, and $\hat{\epsilon}_{i}$ denotes additive Gaussian noise in the frequency domain. The operator $\odot$ represents element-wise multiplication.
@@ 3.3 Homogeneous Reconstruction
[…]
@@ 4 Methods
As shown in Figure 2, CryoFastAR employs a standard vision transformer (ViT) architecture that efficiently encodes particle image features and decodes them into multi-view Fourier planar maps. Specifically, a shared ViT encoder extracts patch-level features, which are then refined through stacked view integration blocks to produce informative representations. These representations are subsequently decoded to Fourier planar maps that encode relative poses (Section 4.1). To effectively train our model on large-scale datasets with high noise levels, we adopt a progressive training strategy, gradually increasing data complexity and noise levels (Section 4.2). During inference, the predicted Fourier planar maps are regressed to explicit 5D poses, facilitating fast ab initio protein reconstruction (Section 4.3).
@@
A fundamental challenge in protein structure recovery is accurately predicting the orientation of a slicing plane in the canonical Fourier space. To address this, we introduce a novel representation termed the Fourier planar map, which encodes per-pixel 3D displacements in Fourier space. These displacements indicate where each 2D Fourier-transformed image lies, in accordance with the Fourier slice theorem. Given a 5D pose $(R,\mathbf{t})$, the Fourier planar map is defined as:
@@
where $X^{0}$ represents a uniformly sampled 2D grid on the plane $z=0$, spanning the range $1,1]^{2}$. Our goal is to develop a neural network that directly takes a set of cryo-EM images as input and predicts the corresponding dense Fourier planar map. Note that the map does not represent actual Fourier coefficients.
@@ 4.1 Model Architecture
Directly applying self-attention across all views introduces quadratic memory and computational complexity, becoming impractical for scenarios involving dozens or hundreds of particle images. To address this, we propose an efficient cross-attention-based approach whose complexity scales linearly with the number of views. Specifically, each decoder layer consists of two key modules: 1) a view integration block, which aggregates target views’ features into a primary reference view (we choose the first view during the training) via cross-attention; and 2) a view updating block, which further refines auxiliary views conditioned on the updated primary features.
@@
Finally, the updated feature representation of each view is decoded into a pixel-wise Fourier planar map accompanied by an auxiliary confidence map, expressed within the reference (first) view’s 3D Fourier coordinate frame. Specifically, two separate downstream heads perform this decoding: one head predicts the reference view’s planar map and the other predicts planar maps for all target views. Formally expressed as:
@@ 4.2 Training
Directly training the model end-to-end on high-noise cryo-EM images poses significant convergence challenges due to the significant differences between cryo-EM data and typical computer vision datasets. To overcome this, we adopt a progressive training approach consisting of three stages. First, we pre-train our model on clean projection images using only two views per particle, facilitating rapid convergence on simplified conditions. Next, we progressively expand training to our full-scale dataset, gradually increasing input views from 2 to 32 and simultaneously reducing the SNR from 10.0 to 0.1, with added CTF distortions. Finally, to mitigate underfitting on experimental data, we fine-tune our model on a small set of real cryo-EM images. This progressive training strategy ensures stable convergence, enabling the model to robustly extract and integrate multi-view features for accurate ab initio reconstruction.
@@ 4.3 Inference
Although initially trained with 32 views, we observed that performance improves with more input views, indicating the model’s ability to integrate multi-view information. During inference, we use 128 views, fixing one as the reference and predicting Fourier planar maps for the remaining views in batches.
Reference View Selection.
Since our model predicts relative poses with respect to a reference view, it is crucial to ensure the reference is of high quality and not a junk particle. We employ a simple yet effective strategy: we sample 64 candidates and select the one with the highest average confidence to other fixed target views as the reference view.
@@ 5 Experiments
To train our model, we construct the first large-scale simulated cryo-EM dataset containing 113,600 atomic structures of protein complexes curated from the Protein Data Bank (PDB). For each atomic structure, we first generate a density volume at a resolution of $128^{3}$ using EMAN2 [42], with voxel sizes automatically determined (ranging from 1Å to 10Å) to fully enclose each structure. From each density volume, we uniformly sample orientations to produce 100 clean 2D projection images at $128\times 128$ resolution. To simulate realistic imaging conditions, we further apply data augmentations, including additive random noise, random in-plane 2D shifts, and CTF modulation. The CTF parameters are sampled from 152,385 CTF parameters in CryoCRAB [5]. For real-world training, we utilize four experimental datasets from CryoDRGN-EMPIAR [53], each containing accurate pre-computed particle poses. For each experimental dataset, we randomly select 30,000 images for training and exclude them from evaluation.
@@ 5.1 Comparison
We compare our approach against three ab initio reconstruction baselines: (1) CryoSPARC, a widely-used software utilizing iterative expectation-maximization and SGD for ab initio reconstruction; (2) CryoSPIN, a neural method directly predicting image poses from single-view inputs; and (3) CryoDRGN2, a hybrid method alternating neural reconstruction and iterative pose search. Notably, unlike these methods, our approach does not require pre-computed CTF parameters for pose estimation. All baseline methods rely on provided CTF parameters during iterative pose refinement or per-scene optimization. Since the official implementation of CryoSPIN does not support shift estimation, we disable random shifts when training it on simulated data and exclude it from comparisons on experimental data. After initial ab initio reconstruction, we further refine our model’s predicted poses using CryoSPARC’s homogeneous refinement to evaluate the quality of our initialization compared to CryoSPARC’s own initialization. Specifically, given our estimated poses, we run CryoSPARC’s reconstruction only followed by the local refinement. For CryoSPARC, we continue to run homogeneous refinement followed by a local refinement.
@@
We compare our method with baseline approaches both qualitatively and quantitatively. As shown in Figure 3, our approach produces reconstructions with superior structural completeness and finer detail preservation. Quantitative results reported in Table 1 further confirm that our method outperforms baselines in pose estimation accuracy and reconstruction quality, achieving significantly lower rotation and translation errors. Remarkably, our method accelerates the reconstruction process by over an order of magnitude (10×), while maintaining or surpassing the quality of all baselines. After refinement, our approach consistently achieves the best overall performance, demonstrating CryoFastAR’s strong initialization quality, computational efficiency, and robustness.
Experimental results. Table 2 summarizes the quantitative results on three experimental cryo-EM datasets including RAG, 50S ribosome, and pre-catalytic Spliceosome [27]. Overall, our approach achieves competitive reconstruction accuracy compared to baseline methods while significantly reducing the computational cost. Specifically, our method achieves approximately a $3.33\times$ speed-up compared to CryoSPARC and over a $33.21\times$ speed-up relative to CryoDRGN2. Notably, our method performs particularly well on the RAG and spliceosome datasets, attaining comparable or better pose estimation accuracy relative to the baselines. However, on the 50S ribosome dataset, our method, although predicting poses in generally correct orientations, yields substantially lower accuracy than the baselines. We hypothesize that this discrepancy arises from the intrinsic structural flexibility and complexity associated with membrane proteins like 50S, posing greater challenges for our feed-forward approach trained primarily on simulated globular structures. As demonstrated in Figure 4, we provide visual comparisons of reconstructions on the experimental spliceosome dataset. These visualizations confirm our quantitative comparisons, demonstrating that our refined reconstruction exhibits improved structural fidelity and better captures the detailed molecular features compared to baseline methods. Other baselines may fail due to lacking the robustness of their heterogeneity without direct supervision during optimization.